# On the optimal prediction of extreme events in heavy-tailed time series with applications to solar flare forecasting

### v1

@@ Abstract
The prediction of extreme events in time series is a fundamental problem arising in many financial, scientific, engineering, and other applications. We begin by establishing a general Neyman-Pearson-type characterization of optimal extreme event predictors in terms of density ratios. This yields new insights and several closed-form optimal extreme event predictors for additive models. These results naturally extend to time series, where we study optimal extreme event prediction for heavy-tailed autoregressive and moving average models. Using a uniform law of large numbers for ergodic time series, we establish the asymptotic optimality of an empirical version of the optimal predictor for autoregressive models. Using multivariate regular variation, we also obtain expressions for the optimal extremal precision in heavy-tailed infinite moving averages, which provide theoretical bounds on the ability to predict extremes in this general class of models. The developed theory and methodology is applied to the important problem of solar flare prediction based on the state-of-the-art GOES satellite flux measurements of the Sun. Our results demonstrate the success and limitations of long-memory autoregressive as well as long-range dependent heavy-tailed FARIMA models for the prediction of extreme solar flares.
@@ 1 Introduction
In many scientific applications, however, one is interested in a particular form of binary prediction. Namely, predicting extreme events $\{Y_{t+h}>y_{0}\}$ associated with the time series $Y$, where $y_{0}$ is a relatively large threshold. An important example motivating our current work is the open problem of forecasting extreme solar flares. Solar flares are sudden, massive eruptions of electromagnetic radiation from the Sun’s atmosphere (NASA Science Editorial Team, 2011; Fletcher et al., 2011). Predicting extreme solar flares is of great interest since they can have adverse effects like radio blackouts (SWPC, n.d.), and they can be accompanied by coronal mass ejections and solar energetic particle events that themselves can have severe effects, such as electrical blackouts (Bolduc, 2002) and irradiation of astronauts (Mertens & Slaba, 2019), respectively. Therefore, extreme solar flare forecasting has been a problem of ongoing interest in the scientific community (see, e.g., Chen et al., 2023, and the reference therein).
@@
Figure 1.1: Left panel: the GOES soft X-ray flux time series over the window 9/1/2011-5/31/2014. Right panel: the time series between 11:30 PM on 3/6/2014 and 1:00 AM on 3/7/2014, during which a strong solar flare occurred.
@@
The strength of a solar flare is defined as the peak value of the soft solar X-ray flux while it is ongoing (Fletcher et al., 2011). The left panel of Figure 1.1 displays a time series of flux measurements from the period 9/1/2011-5/31/2014 obtained from the Geostationary Operational Environmental Satellites (GOES) (cf Section 5); the right panel shows the same time series, but around the time of a strong solar flare that occurred on the night of 3/6/2014-3/7/2014. The key features of the flux time series are heavy tails and complex temporal dependence (Stanislavsky et al., 2009). The main objective is to provide timely forecasts of extreme flaring events $\{Y_{t+h}>y_{0}\}$, where $Y_{t+h}$ is the flux at a future time $t+h$ and $y_{0}$ is a high threshold. This is one formulation of the so-called operational flare forecasting problem. A plethora of modern machine learning tools have been used for operational flare forecasting, but their accuracy has been far from satisfactory; though they have had tremendous success in other prediction tasks, they have not managed to surpass more basic approaches, and the problem is still largely open (Barnes et al., 2016; Leka et al., 2019a, b).
@@
In this and many similar contexts, a natural question arises as to what are the optimal predictors and the ultimate theoretical limits of accuracy in predicting extreme events in time series, regardless of the methodology employed. In this paper, we address the fundamental problem of optimal prediction of extreme events $\{Y_{t+h}>y_{0}\}$ in time series. Motivated by the solar flare prediction problem, we focus on heavy-tailed time series models though many of our results apply in greater generality. Somewhat surprisingly, this important binary prediction problem has been studied much less than other prediction problems in the literature. Therefore, in Section 2, we begin by outlining a framework for the optimal prediction of extreme events. Briefly, suppose that the target event occurs at a rate $1-p=\mathbb{P}[Y_{t+h}>y_{0}]$, where $p\approx 1$. Then, all predictors of these events can be written in the form $\{\widehat{Y}_{t}(h)>\widehat{y}_{0}\}$, where $\widehat{Y}_{t}(h)=g(Y_{s},\ s\leq t)$ is some statistic of the observed data. We say that the predictor is calibrated if it raises alarms at a specified rate, i.e., $\mathbb{P}[\widehat{Y}_{t}(h)>\widehat{y}_{0}]=1-q$ for a specified $q$; typically, $q=p$. The precision of a calibrated predictor is then
${\rm prec}(\widehat{Y}_{t}(h)):=\mathbb{P}[Y_{t+h}>y_{0}|\widehat{Y}_{t}(h)>\widehat{y}_{0}].$ (1.1)
Naturally, a predictor will be said to be optimal if it maximizes the above precision among all calibrated predictors. Thus, calibrated optimal predictors are ones that maximize the precision while controlling the overall alarm rate (see also Sections 2.1 and A.1.3 for optimality under other modes of calibration).
@@
In Section 2.3, motivated by our main objective to predict extreme events, we study the asymptotic aspects of optimal prediction as the threshold $y_{0}$ grows, or equivalently, as $p\uparrow 1$. It turns out that the limiting precision in (1.1) is in fact the upper tail dependence coefficient between $Y_{t+h}$ and $\widehat{Y}_{t}(h)$. Tail dependence has been widely used and studied in extreme value theory (see, e.g., Russell et al., 2016; Janßen et al., 2023, and the references therein). However, to the best of our knowledge, our work is the first to explicitly characterize the optimal extreme event predictors, namely, the ones that lead to the strongest tail dependence (see, Theorem 2.3, below). This naturally leads us to establish new results on the optimal extremal precision for prediction of extreme events in heavy-tailed linear time series (cf Section 3.3). These results characterize the best possible asymptotic precision that can be achieved by any calibrated predictor of the target extreme events. Interestingly, for a large class of causal infinite moving average models with absolutely monotone sequence of coefficients, optimal extreme event predictors are simply achieved by using the most recent observation (Proposition 3.2). In general, however, the optimal predictors for moving average models involve the complete history of the time series. In the important special case of autoregressive models, these predictors can be computed with a finite sample (see Section 3.1). In Section 3.2, we propose a practical methodology for the estimation of the optimal predictors in autoregressive time series. Using results on uniform large numbers for empirical processes of ergodic time series (Adams & Nobel, 2012), we establish the asymptotic calibration and optimality of the predictors.
@@
In Section 4, we investigate the finite-sample performance of our optimal event prediction methodology using simulated data (Figure 4.1). In Section 5, we present an extensive application of the optimal extreme event prediction methodology to the state-of-the-art GOES flux time series data. We model the data with both autoregressive and heavy-tailed long-range dependent (FARIMA) time series models, as advocated for in works such as Burnecki & Weron (2014). The optimal extreme event predictors under these models provide new insights on the challenging solar flare prediction problem. Specifically, they indicate that both heavy tails and long-range dependence should be accounted for when predicting extreme solar flares. The resulting empirical precisions and true skill statistics (Table 5.1) provide lower bounds on the optimal prediction accuracy one can hope to achieve when predicting solar flares using time series methods applied to historical GOES flux data. Operational flare forecasting, however, may be significantly improved through the use of covariate information such as images of the Sun at various wavelengths. Inference for optimal extreme event predictors in this context is a future direction of research, which we plan to pursue with the guidance of the general characterization results in Theorems 2.1 and 2.3.
@@
The rest of the paper is structured as follows. In Section 2.1, we give a framework and general characterization results for optimal event prediction. We study the optimal extremal precision in Section 2.3. Section 3.1 applies and extends these results to the important class of linear heavy-tailed time series models. Section 3.2 develops statistical inference for the optimal predictors in autoregressive time series, while Section 3.3 establishes general theoretical results on the optimal extremal precision in heavy-tailed moving average models. Section 4 studies optimal extreme event predictors using simulated data, while Section 5 applies prediction methods to the solar flare prediction problem using GOES flux data from the most active periods of two recent solar cycles. Technical proofs and further details are given in the Appendix.
@@ 2.1 A General Characterization of Optimal Predictors
We will refer to $F_{\xi}^{\leftarrow}(p)$ as to the $p$th quantile of $\xi$, for $p\in(0,1)$. We shall need the following result. Its proof is given in Section A.1.1.
@@
The notion of precision in Definition 2.1 is closely related to the notion of precision arising in the evaluation of classifiers. See Fawcett (2006) and Section A.1.3.
@@
In applications, especially when predicting extreme events $(p\approx 1)$, we often consider the balanced case where
$p=q.$
This requirement is natural as it mandates that alarms be raised only as frequently as events occur. However, a practitioner may want to construct predictors for varying levels of the alarm rate $1-q$. This naturally leads to the notion of ROC curve optimality of families of predictors that differ in their alarm rate. We discuss this notion in Section A.1.3.
@@ 2.3 Asymptotic Aspects of Optimal Prediction: Extremal Precision
In view of Example 2.2, without loss of generality we may assume that $(Y,X)$ is Gaussian. Recalling Example 2.1, we can write $Y=h(X)+\epsilon,$ where $h(X):=a^{\top}X$ and $\epsilon$ are independent Gaussian random variables and hence $h(X)=a^{\top}X$ yields the optimal predictor for the events $\{Y>F_{Y}^{\leftarrow}(p)\}$ (Theorem 2.2). Thus, by Theorem 2.3,
$\lambda^{(opt)}(Y,X)=\lambda(Y,a^{\top}X).$
However, since $(Y,X)$ has a non-singular covariance matrix, ${\rm Corr}(Y,a^{\top}X)<1$ and the tail-dependence coefficient $\lambda(Y,a^{\top}X)=0$. See, for example, Relation (5.1) in Ledford & Tawn (1996) or Table 4.1 in Joe (2015). ∎
@@ 3.1 Explicit Optimal Predictors
We shall assume that the innovations satisfy Assumption 3.1 and that for each $t$, $\epsilon_{t}$ is independent from the history of the process up to time $t-1$, i.e., the $\sigma$-field ${\cal H}_{t-1}:=\sigma\{Y_{s},\ s\leq t-1\}$. Under the usual conditions that the complex polynomial $\pi(z):=1-\sum_{i=1}^{d}\phi_{i}z^{i}$ has its roots outside the unit circle, we have that the autoregressive equation (3.5) has a unique stationary solution:
@@
where $\sum_{i=0}^{\infty}a_{i}z^{i}=1/\pi(z),\ |z|\leq 1$ is the series representation of the function $1/\pi(z)$, which is analytic in the unit circle $\{|z|\leq 1\}$. The coefficients $a_{i}$ in the moving average representation (3.6) decay exponentially fast and therefore, since $\mathbb{E}[|\epsilon_{i}|^{\delta}]<\infty$ (Assumption 3.1), the series in (3.6) is convergent almost surely (see, e.g., Propositions 13.3.1 and 13.3.2 in Brockwell & Davis, 1991).
The following lemma will help us derive the optimal predictor for the model in (3.5). For $t\geq u$, we use the notation $Y_{t:u}$ to succinctly represent the vector $(Y_{t},\ldots,Y_{u})^{\top}$. The proof is given in Section A.2.1.
@@ 3.2 Statistical Inference for Optimal Predictors in AR($d$) Models
The optimal predictor for AR$(d)$ models requires knowledge of $\phi$, as well as the distribution of $\widehat{Y}_{t+h}^{\text{(opt)}}$ in order to calibrate the predictor with the appropriate quantile. In practice, both $\phi$ and the quantile need to be estimated. In this section, we study a natural plug-in approximation to the optimal predictor $\widehat{Y}_{t+h}^{\text{(opt)}}$ and its calibration from training data. We establish the asymptotic consistency and optimality of the inferential procedure as the size of the training sample goes to infinity. We do so under very general conditions using a uniform law of large numbers for empirical processes established in Adams & Nobel (2012).

### v2

@@ Abstract
The prediction of extreme events in time series is a fundamental problem arising in many financial, scientific, engineering, and other applications. We begin by establishing a general Neyman-Pearson-type characterization of optimal extreme event predictors in terms of density ratios. This yields new insights and several closed-form optimal extreme event predictors for additive models. These results naturally extend to time series, where we study optimal extreme event prediction for both light- and heavy-tailed autoregressive and moving average models. Using a uniform law of large numbers for ergodic time series, we establish the asymptotic optimality of an empirical version of the optimal predictor for autoregressive models. Using multivariate regular variation, we obtain an expression for the optimal extremal precision in heavy-tailed infinite moving averages, which provides theoretical bounds on the ability to predict extremes in this general class of models. We address the important problem of predicting solar flares by applying our theory and methodology to a state-of-the-art time series consisting of solar soft X-ray flux measurements. Our results demonstrate the success and limitations in solar flare forecasting of long-memory autoregressive models and long-range-dependent, heavy-tailed FARIMA models.
@@ 1 Introduction
In many scientific applications, however, one is interested in a particular form of binary prediction. Namely, predicting extreme events $\{Y_{t+h}>y_{0}\}$ associated with the time series $Y$, where $y_{0}$ is a relatively large threshold. An important example motivating our current work is the problem of forecasting extreme solar flares, which has long been a problem of interest in the scientific community (see, e.g., Chen et al., 2023, and the references therein). Solar flares are sudden, massive eruptions of electromagnetic radiation from the Sun’s atmosphere (NASA Science Editorial Team, 2011; Fletcher et al., 2011). Predicting extreme solar flares is of great interest since they can have adverse effects like radio blackouts (SWPC, n.d.b), and they can be accompanied by coronal mass ejections (CMEs) and solar energetic particle events that themselves can have severe effects, such as electrical blackouts (Bolduc, 2002) and irradiation of astronauts (Mertens & Slaba, 2019), respectively. Notably, in February 2022, a CME associated with a strong solar flare resulted in the destruction of 38 Starlink satellites (Kataoka et al., 2022).
@@
Figure 1.1: Left: the GOES soft X-ray flux time series over the window 1 September 2011 to 31 May 2014. Right: the time series between 11:30 PM on 6 March 2014 and 1:00 AM on 7 March 2014, during which a strong solar flare occurred.
@@
The strength of a solar flare is defined as the peak value of the soft solar X-ray flux while it is ongoing (Fletcher et al., 2011). The left panel of Figure 1.1 displays a time series of flux measurements from the period 1 September 2011 to 31 May 2014 obtained from the Geostationary Operational Environmental Satellites (GOES) (cf. Section 4); the right panel shows the same time series, but around the time of a strong solar flare that occurred on the night of 6–7 March 2014. The key features of the flux time series are heavy tails and complex temporal dependence (Stanislavsky et al., 2009). The main objective is to provide timely forecasts of extreme flaring events $\{Y_{t+h}>y_{0}\}$, where $Y_{t+h}$ is the flux at a future time $t+h$ and $y_{0}$ is a high threshold. This is one formulation of the so-called operational flare forecasting problem. A plethora of modern machine learning tools have been used for operational flare forecasting, but their accuracy has been far from satisfactory; though they have had tremendous success in other prediction tasks, they have not managed to surpass more basic approaches, and the problem is still largely open (Barnes et al., 2016; Leka et al., 2019a, b). Note that our framing of the flare forecasting problem can be applied to other time series forecasting problems; for example, it can be applied to the problem of forecasting the solar radio flux, an important proxy for solar activity that has been widely studied (SWPC, n.d.a; Gromenko et al., 2017).
@@
In the flare forecasting context and similar contexts, a natural question arises as to what are the optimal predictors and the ultimate theoretical limits of accuracy in predicting extreme events in time series, regardless of the methodology employed. In this paper, we address the fundamental problem of optimal prediction of extreme events $\{Y_{t+h}>y_{0}\}$ in time series. Motivated by the solar flare prediction problem, we focus on heavy-tailed time series models though many of our results apply in greater generality. Somewhat surprisingly, this important binary prediction problem has been studied much less than other prediction problems in the literature. Therefore, in Section 2, we begin by outlining a framework for the optimal prediction of extreme events. Briefly, suppose that the target event occurs at a rate $1-p=\mathbb{P}[Y_{t+h}>y_{0}]$, where $p\approx 1$. Then, all predictors of these events can be written in the form $\{\widehat{Y}_{t}(h)>\widehat{y}_{0}\}$, where $\widehat{Y}_{t}(h)=g(Y_{s},\ s\leq t)$ is some statistic of the observed data. We say that the predictor is calibrated if it raises alarms at a specified rate, i.e., $\mathbb{P}[\widehat{Y}_{t}(h)>\widehat{y}_{0}]=1-q$ for a specified $q$; typically, $q=p$. The precision of a calibrated predictor $\{\widehat{Y}_{t}(h)>\widehat{y}_{0}\}$ is
$\text{prec}_{Y_{t+h}>y_{0}}(\widehat{Y}_{t}(h)>\widehat{y}_{0}):=\mathbb{P}[Y_{t+h}>y_{0}|\widehat{Y}_{t}(h)>\widehat{y}_{0}].$ (1.1)
Note that if the current time is $t$, then the event $\{Y_{t+h}>y_{0}\}$ is uncertain; it may or may not occur. If we raise an alarm at time $t$, we quantify our uncertainty about whether $Y_{t+h}$ will be extreme using the precision. We give a more general definition of our notion of precision in Definition 2.1. It is the population analogue of the sample precision, a metric that is widely used in the evaluation of classifiers and is defined as the proportion of alarms that were followed by the event of interest (Fawcett, 2006). See Remark A.5 for more detail on the motivation for the definition. A predictor will be said to be optimal if it maximizes the precision among all calibrated predictors. Thus, calibrated optimal predictors are ones that maximize the precision while controlling the alarm rate (see also Sections 2.1 and A.1.3 for optimality under other modes of calibration).
@@
In Section 2.3, motivated by our main objective to predict extreme events, we study the asymptotic aspects of optimal prediction as the threshold $y_{0}$ grows, or equivalently, as $p\uparrow 1$. It turns out that the limiting precision in (1.1) is in fact the upper tail dependence coefficient between $Y_{t+h}$ and $\widehat{Y}_{t}(h)$. Tail dependence has been widely used and studied in extreme value theory (see, e.g., Russell et al., 2016; Janßen et al., 2023, and the references therein). However, to the best of our knowledge, our work is the first to explicitly characterize the optimal extreme event predictors, namely, the ones that lead to the strongest tail dependence (see Theorem 2.3, below). This naturally leads us to establish new results on the optimal extremal precision for prediction of extreme events in heavy-tailed linear time series (cf. Section 3.3, particularly Theorem 3.4). These results characterize the best possible asymptotic precision that can be achieved by any calibrated predictor of the target extreme events. Interestingly, for a large class of causal infinite moving average models with absolutely monotone sequence of coefficients, optimal extreme event predictors are simply achieved by using the most recent observation (Proposition 3.2). In general, however, the optimal predictors for moving average models involve the complete history of the time series. In the important special case of autoregressive models, these predictors can be computed with a finite sample (see Section 3.1). In Section 3.2, we propose a practical methodology for the estimation of the optimal predictors in autoregressive time series. Using results on uniform large numbers for empirical processes of ergodic time series (Adams & Nobel, 2012), we establish the asymptotic calibration and optimality of the predictors. Table 1.1 summarizes the results we present for the kinds of time series models we consider in this paper, infinite moving average and autoregressive models.
Model Oracle Estimator Inference Extremal Oracle Precision
AR($d$) Theorem 3.2 Section 3.2.1 Theorem 3.3 Numerical
MA($\infty$) Theorem 3.1 Not Included Open Problem Theorem 3.4
FARIMA$(0,d,0)$ Theorem 3.1 Not Included Open Problem Theorem 3.4
Table 1.1: A summary of the results we present for the models we consider in this paper. Since FARIMA$(0,d,0)$ models are MA($\infty$)∞ ) models, the results given for the former are the same as those given for the latter. For those two kinds of models, estimators that approximate the optimal or oracle predictor can be obtained by truncating the oracle predictor; they are not included in this paper. The inferential properties of these estimators remain to be studied. The extremal precision of the oracle estimator for AR($d$) models can be obtained by computing the MA($\infty$)∞ ) representation and then applying Theorem 3.4.
@@
In Section 4, we present an extensive application of the optimal extreme event prediction methodology to the state-of-the-art GOES flux time series data. We model the data with both autoregressive and heavy-tailed long-range dependent (FARIMA) time series models, as advocated for in works such as Burnecki & Weron (2014). The optimal extreme event predictors under these models provide new insights on the challenging solar flare prediction problem. Specifically, they indicate that both heavy tails and long-range dependence should be accounted for when predicting extreme solar flares. The resulting empirical precisions and true skill statistics (Table 4.1) provide lower bounds on the optimal prediction accuracy one can hope to achieve when predicting solar flares using time series methods applied to historical GOES flux data. Operational flare forecasting, however, may be significantly improved through the use of covariate information such as images of the Sun at various wavelengths. Inference for optimal extreme event predictors in this context is a future direction of research, which we plan to pursue with the guidance of the general characterization results in Theorems 2.1 and 2.3.
@@
The rest of the paper is structured as follows. In Section 2.1, we give a framework and general characterization results for optimal event prediction. We study the optimal extremal precision in Section 2.3. Section 3.1 applies and extends these results to the important class of linear heavy-tailed time series models. Section 3.2 develops statistical inference for the optimal predictors in autoregressive time series, while Section 3.3 establishes general theoretical results on the optimal extremal precision in heavy-tailed moving average models. Section 4 applies prediction methods to the solar flare prediction problem using GOES flux data from the most active periods of two recent solar cycles. Technical proofs and further details are given in the Appendix.
@@ 2.1 A General Characterization of Optimal Predictors
We will refer to $F_{\xi}^{\leftarrow}(p)$ as to the $p$th quantile of $\xi$, for $p\in(0,1)$. We shall need the following result.
@@
The notion of precision in Definition 2.1 is closely related to the notion of precision arising in the evaluation of classifiers. See Fawcett (2006) and Remark A.5.
@@
In applications, especially when predicting extreme events $(p\approx 1)$, we often consider the balanced case where $p=q$. This requirement is natural as it mandates that alarms be raised only as frequently as events occur. However, a practitioner may want to construct predictors for varying levels of the alarm rate $1-q$. This naturally leads to the notion of ROC curve optimality of families of predictors that differ in their alarm rate. We discuss this notion in Section A.1.3.
@@ 2.3 Asymptotic Aspects of Optimal Prediction: Extremal Precision
In view of Example 2.2, without loss of generality we may assume that $(Y,X)$ is Gaussian. Recalling Example 2.1, we can write $Y=h(X)+\epsilon,$ where $h(X):=a^{\top}X$ and $\epsilon$ are independent Gaussian random variables and hence $h(X)=a^{\top}X$ yields the optimal predictor for the events $\{Y>F_{Y}^{\leftarrow}(p)\}$ (Theorem 2.2). Thus, by Theorem 2.3, $\lambda^{(opt)}(Y,X)=\lambda(Y,a^{\top}X)$. However, since $(Y,X)$ has a non-singular covariance matrix, ${\rm Corr}(Y,a^{\top}X)<1$ and the tail-dependence coefficient $\lambda(Y,a^{\top}X)=0$. See, for example, Relation (5.1) in Ledford & Tawn (1996) or Table 4.1 in Joe (2015). ∎
@@ 3.1 Explicit Optimal Predictors
We shall assume that the innovations satisfy Assumption 3.1 and that for each $t$, $\epsilon_{t}$ is independent from the history of the process up to time $t-1$, i.e., the $\sigma$-field ${\cal H}_{t-1}:=\sigma\{Y_{s},\ s\leq t-1\}$. Under the usual conditions that the complex polynomial $\pi(z):=1-\sum_{i=1}^{d}\phi_{i}z^{i}$ has its roots outside the unit circle, the autoregressive equation (3.5) has a unique stationary solution:
@@
where $\sum_{i=0}^{\infty}a_{i}z^{i}=1/\pi(z),\ |z|\leq 1$ is the series representation of the function $1/\pi(z)$, which is analytic in the unit circle $\{|z|\leq 1\}$. The coefficients $a_{i}$ in the moving average representation (3.6) decay exponentially fast and therefore, since $\mathbb{E}[|\epsilon_{t}|^{\delta}]<\infty$ (Assumption 3.1), the series in (3.6) is convergent almost surely (see, e.g., Propositions 13.3.1 and 13.3.2 in Brockwell & Davis, 1991).
The following lemma will help us derive the optimal predictor for the model in (3.5). For $t\geq u$, we use the notation $Y_{t:u}$ to succinctly represent the vector $(Y_{t},\ldots,Y_{u})^{\top}$.
@@ 3.2 Statistical Inference for Optimal Predictors in AR($d$) Models
The optimal predictor for AR$(d)$ models requires knowledge of $\phi$, as well as the distribution of $\widehat{Y}_{t+h}^{\text{(opt)}}$ in order to calibrate the predictor with the appropriate quantile. In practice, both $\phi$ and the quantile need to be estimated. In this section, we study a natural plug-in approximation to the optimal predictor $\widehat{Y}_{t+h}^{\text{(opt)}}$ and its calibration from training data. We establish the asymptotic consistency and optimality of the inferential procedure as the size of the training sample goes to infinity. We do so under very general conditions using a uniform law of large numbers for empirical processes established in Adams & Nobel (2012) (see the proof of Proposition 3.1).